OpenStudy (anonymous):

Write y = 1/6x + 5 in standard form using integers. A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5 A, B, C, or D? Explain, please.

hartnn (hartnn):

Multiply 6 on both sides of y = 1/6x + 5 .

OpenStudy (anonymous):

So multiply 1/6 and 5?

hartnn (hartnn):

no, multiply 6. to each term 6*y=? 6(1/6) x =? 6*5=?

OpenStudy (anonymous):

Okay, and how would I multiply 6(1/6)?

hartnn (hartnn):

how do u think ? just give it a try ...

OpenStudy (anonymous):

6*y= I have no idea because y is a letter.. Lol perhaps some help on that bit? I think I might have to switth 1/6 to a decimal in order to multiply it in this situation, so I'm gonna go with .16, and .16*6=.96.. And 6*5 = 30

OpenStudy (anonymous):

switch*

hartnn (hartnn):

OMG! 6*(1/6) = 6/6 = 1. so u have 6y=x+30 subtract x from both sides, what u get ?

OpenStudy (anonymous):

How do I subtract x from both sides?

hartnn (hartnn):

-x+6y = x-x +30 what is x-x = ?

OpenStudy (anonymous):

I am unsure.

hartnn (hartnn):

what is 1-1=?

OpenStudy (anonymous):

0

hartnn (hartnn):

so x-x=0 finally, u have -x+6y=30 got this ?

OpenStudy (anonymous):

Yes, I think so. Thank you.

hartnn (hartnn):

welcome ^_^

OpenStudy (lgbasallote):

just so you know...the decimal of 1/6 is 0.17 not 0.16

OpenStudy (anonymous):

Ah, Thank you (: